What is the velocity of the wave in a guitar string having a length of 0.65m and a mass of 3.5g when subjected to a tension of 250N?

1
Expert's answer
2022-01-24T11:09:34-0500

Given:

"L=0.65\:\rm m"

"m=3.5*10^{-3}\:\rm kg"

"T=250\:\rm N"


The velocity of the wave in a guitar string

"v=\sqrt{\frac{T}{m/L}}=\sqrt{\frac{250}{3.5*10^{-3}/0.65}}=331\:\rm m/s"